OpenStudy (anonymous):

if x^2/36 +y^2/16 and y(1)=3.94405 find y'(1) by implicit differentiation.

OpenStudy (anonymous):

sorry, if x^2/36 +y^2/16 = something?

OpenStudy (anonymous):

opps if x^2/36 +y^2/16 =1

OpenStudy (anonymous):

ok. so x/13+yy'/8=0 plug in 1 for x and 3.94405 for y 1/13+(3.94405/8)y'=0 y'(1)=-(1/13)(8/3.94405)

OpenStudy (australopithecus):

2x/36 + (2y/16)y' = 0 next sub in values -2x(16)/36(2y) = y' -2(1)(16)/( 36(2)(3.94405) ) = y'

OpenStudy (anonymous):

so what will the answer equal?

OpenStudy (australopithecus):

There is your answer, I hope you understand this material

OpenStudy (anonymous):

thank you, and yes im starting to get the hang of it